# Planning-Aware Code Infilling via Horizon-Length Prediction

### v3

@@ Abstract
Yifeng Ding1††thanks: Work done during a research internship at AWS AI Labs. Hantian Ding2 Shiqi Wang2 Qing Sun2 Varun Kumar2 Zijian Wang2 University of Illinois Urbana-Champaign1 AWS AI Labs2 yifeng6@illinois.edu {dhantian,zijwan}@amazon.com
@@ 1 Introduction
Fill-in-the-Middle (FIM), or infilling, has become essential for modern code development, where programmers frequently need to insert or modify code between existing sections rather than writing linearly from start to end (Bavarian et al., 2022; Fried et al., 2023). While large language models have shown remarkable capabilities in code generation, the FIM task poses unique challenges that go beyond traditional left-to-right generation. A model must not only generate code that follows from the preceding context (), but also smoothly connect to the subsequent code () – a task that requires careful planning and foresight.
Current approaches to FIM typically reorder the input sequence and apply standard next-token prediction (NTP) training (Lozhkov et al., 2024; Guo et al., 2024; DeepSeek-AI et al., 2024; Hui et al., 2024). However, as illustrated in Figure 1, this methodology has a fundamental limitation: models struggle to maintain coherence over longer sequences and often fail to create smooth transitions to the right context (Bavarian et al., 2022). While existing benchmarks attempt to address this through rule-based post-processing (e.g., truncating generated code based on line counts (Zhang et al., 2023; Wu et al., 2024) or program structure (Ding et al., 2023; Gong et al., 2024)), such methods rely on dataset-specific assumptions that do not generalize to real-world scenarios where both left and right contexts can be arbitrary.
@@
| RepoEval (Zhang et al., 2023) | Truncate generation to the same number of lines as in ground truth. |
| CrossCodeLongEval (Wu et al., 2024) | |
| CrossCodeEval (Ding et al., 2023) | Truncate generation at the first complete statement. |
| SAFIM (Gong et al., 2024) | Stop when the target program structure in ground truth is generated. |
@@ 2 Post-processing for Fill-in-the-Middle
Most existing FIM works rely on post-processing to truncate code completions generated by LLMs for infilling tasks (Gong et al., 2024; Zhang et al., 2023; Ding et al., 2023; Wu et al., 2024). While such post-processing can enhance the FIM performance, we argue that they fundamentally depend on dataset-specific assumptions that make them impractical for real-world scenarios (§2.1). Through evaluation, we show that the performance of FIM existing code models drops significantly without post-processing, suggesting that post-processing conceals the fundamental struggles of models with code filling (§2.2). Furthermore, we show that this limitation stems from models’ inability to plan coherent completions given a fixed – a challenge that post-processing alone cannot address.
@@ 3 Horizon-Length Prediction
Next-token prediction (NTP) training is conducted on the document , whose goal is to minimize the following cross-entropy loss (where refers to LLMs being trained):
@@ 4 Experiments
[…]
Training. We conduct continual pre-training on a set of code LLMs of different model families and sizes to validate the effectiveness of HLP. Specifically, DeepSeek-Coder-Base 1.3B/6.7B (Guo et al., 2024) and StarCoder2 3B/7B (Lozhkov et al., 2024) are involved in our experiments. We use AdamW (Loshchilov and Hutter, 2019) as the optimizer with and . We use a cosine learning rate scheduler with a peak learning rate equal to that at the pre-training end. All models are trained for 200K steps with a global batch size of 512.
Dataset. We use a subset of the-stack-v2 (Lozhkov et al., 2024) including Python, Java, C++, and C#. Following existing works (Guo et al., 2024; Lozhkov et al., 2024), FIM rate is set to . We employ Best-fit Packing (Ding et al., 2024) to group multiple files into each training sequence while masking out cross-file attention. The -- split is applied to each file independently rather than the whole training sequence.
@@ 4.1 Syntax-Aware Multilingual Code FIM
We use SAFIM (Gong et al., 2024), a syntax-aware and multilingual code Fill-in-the-Middle benchmark, to evaluate the effectiveness of HLP. SAFIM focuses on syntax-aware completions of program structures, covering algorithmic block (i.e., Algo and Algo), control-flow expression (i.e., Control), and API function call (i.e., API). It consists of 17,720 examples from four different programming languages, including Python, Java, C++, and C#. SAFIM employs execution-based evaluation and reports pass@1 as the evaluation metric. As shown in Table 3, compared with NTP only, adding HLP achieves up to 5% improvements on average across all the studied code LLMs. Importantly, the improvement is consistent across languages and program structures, demonstrating HLP’s ability to enhance FIM capabilities regardless of language or completion context.
@@ 4.2 Repository-Level Cross-File Code FIM
In addition to single-file FIM evaluation with SAFIM, we also evaluate the effectiveness of HLP on repository-level code Fill-in-the-Middle in cross-file scenarios via CrossCodeEval (Ding et al., 2023), CrossCodeLongEval (Wu et al., 2024), and RepoEval (Zhang et al., 2023). CrossCodeEval (Python) and CrossCodeLongEval are two repository-level cross-file benchmarks that leverage more than 1500 raw Python repositories to construct 12,665 examples across line, chunk, and function completion tasks, which are used for a more rigorous evaluation. RepoEval is another repository-level cross-file code completion benchmark consisting of 3,655 line, API, and function completion tasks created from 32 Python repositories. We follow existing work (Wu et al., 2024) to evaluate the model’s FIM performance on these benchmarks and use Exact Match (EM) and Edit Similarity (ES) as our evaluation metrics. As shown in Table 4, adding HLP provides consistent improvements for all models across different benchmarks and completion tasks. Specifically, HLP achieves up to 24% improvements on EM and 9% improvements on ES relatively, showing its significant effectiveness.
@@ 4.3 Code Repair via Fill-in-the-Middle
To assess HLP’s impact on practical applications beyond code completion, we evaluate the performance of code repair using Defects4J (Just et al., 2014). Defects4J consists of open-source bugs found across 15 Java repositories. Following existing works (Xia et al., 2023; Xia and Zhang, 2023), we collect 313 single-hunk bugs from Defects4J that can be fixed by replacing or adding a continuous code hunk. Specifically, for each bug, models are prompted to generate the correct code hunk (i.e., patch) given the left and right contexts of the buggy code hunk, with correctness verified by project test suites. As shown in the “Code Repair” section of Table 5, adding HLP during training results in relatively up to 18% more bugs fixed by the model222Note that Defects4J is a small dataset with only 313 examples and models can only solve 30-60 out of those, which makes it hard to obtain statistically significant differences with greedy decoding., suggesting that enhanced planning capabilities translate directly to better bug-fixing performance.
@@ 4.4 Code Reasoning via Fill-in-the-Middle
We further examine whether HLP’s planning benefits extend to code reasoning tasks. We use CRUXEval (Gu et al., 2024) which comprises 800 Python functions paired with two distinct tasks: CRUXEval-I, where LLMs need to predict the input from the known output, and CRUXEval-O, where LLMs are required to predict the output based on the given input.
@@ 5.1 NTP Alone Cannot Yield Horizon Awareness
To quantify this, we design a probing task that attempts to predict the remaining completion length from s’ hidden states. We extract hidden states from 20K code snippets (approximately 7.8M tokens) using models trained with and without HLP and split the data to ensure no sequence-level overlap between train and test. Using these hidden states as input and normalized remaining token counts as targets, we fit linear regression models while keeping the underlying transformer parameters frozen.
@@ 6 Related Work
Fill-in-the-Middle for Code Language Models Large language models trained on massive source code data have demonstrated great potential in various applications for software development. While early models such as Codex (Chen et al., 2021) and CodeGen (Nijkamp et al., 2023) only support Left-to-Right (L2R) generation, Fill-in-the-Middle (or infilling) has attracted increased attention because right context naturally carries an indispensable part of information for completing code in the middle (Fried et al., 2023; Bavarian et al., 2022). Subsequently, FIM training has become a common practice widely adopted by most code LLMs, such as StarCoder (Li et al., 2023; Lozhkov et al., 2024), DeepSeek-Coder (Guo et al., 2024; DeepSeek-AI et al., 2024), and Code Llama (Rozière et al., 2023).
@@
An alternative approach is to train two language models in different directions, with one generating from left to right and the other from right to left, and have the two generations meet in the middle (Nguyen et al., 2023). Nevertheless, the L2R model does not have access to the right context, and vice versa, which impedes holistic planning that takes into account the context from both sides.
Planning and Lookahead in Language Generation Standard decoder-only models are trained with next-token prediction and used to sequentially predict one token at a time, conditioned only on past tokens, in an auto-regressive manner. One drawback of this paradigm is that models are not aware of future tokens during decoding. The token that maximizes the conditional probability at current step may lead to suboptimal continuation, and consequently the model can fail to compose a fluent and sensible generation that meets human requirements. Various decoding techniques have been proposed to address the problem through tree search with lookahead heuristics, particularly for constrained generation problems (Lu et al., 2022; Huang et al., 2024). While these methods are training-free, they inevitably incur additional cost of inference complexity.
Apart from those, Gloeckle et al. (2024) proposed to predict multiple tokens from a single hidden state during both training and inference, which was shown to achieve stronger performance on coding tasks with no computation overhead. While multi-token prediction enhances models’ planning capability within the tokens predicted together (), we argue that with a small , the limited horizon is usually insufficient for planning in the case of infilling as the connection from to only happens towards the end of the generation. In contrast, HLP adopts a global and arbitrary long horizon over all future tokens by counting the remaining generation budget, which more effectively helps models to close the generation fluently with early planning444Please refer to Appendix §A.3 for more details..
@@ A.3 Comparing HLP with Multi-token Prediction
[…]
As discussed in §6, we argue that multi-token prediction (Gloeckle et al., 2024) is insufficient for planning in FIM, because multi-token prediction only enhances models’ planning capability for a short and limited horizon, which does not suites FIM well as the connection from to happens over a long horizon. Instead, HLP focuses on long-horizon planning and is more effective for FIM. We conduct an experiment following the same settings in §4 to compare HLP with multi-token prediction on DeepSeek-Coder-Base 1.3B by predicting the next 4 tokens (Gloeckle et al., 2024). We report their performance on SAFIM. As shown in Table 10, while adding HLP to NTP largely improves the model’s performance on SAFIM, multi-token prediction fails to do so. These results provide empirical evidence that long-horizon planning capabilities brought by HLP is essential for advancing FIM performance.
@@ A.4 HLP Improves Self-Infilling Performance
Recently, some strategies have been proposed to conduct L2R code generation based on FIM capabilities and self-infilling (Zheng et al., 2024) is one of the representative strategies of this kind. Given an initial input prompt , self-infilling performs L2R code generation by using FIM decoding to (1) generate based on and then (2) generate based on and previously generated . Furthermore, self-infilling proposes a looping mechanism to improve the generated code iteratively, where it first uses L2R decoding to generate a new ′ based on and previously generated and then uses FIM decoding to generate a new ′ based on and this newly generated ′. This looping procedure can be continued for multiple rounds to obtain greater improvements.
@@
[…]
Since self-infilling relies on FIM capabilities for L2R code generation, it is interesting to study whether HLP can also enhance its performance. To this end, we evaluate the self-infilling performance of DeepSeek-Coder-Base 1.3B trained with and without HLP on HumanEval using greedy decoding, following the setting of the original paper (Zheng et al., 2024). We report models’ performance on HumanEval with ranging from 0 to 4, where denotes the number of times the decoding process goes through the loop and represents that the looping mechanism is not activated. As shown in Table 11, the model trained with HLP consistently outperforms the model trained without HLP across different number of self-improving steps. Furthermore, with more self-improving steps, while the performance of the model trained without HLP gets stuck, the performance of the model trained with HLP continues to showcase steady improvements.
@@ A.5 Effect of HLP on Left-to-Right Performance
While HLP have significantly improved the FIM performance of LLMs, we also study its impact on the L2R code completion. To this end, we evaluate L2R performance on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) with DeepSeek-Coder-Base 1.3B. We further employ HumanEval+ and MBPP+ from EvalPlus Liu et al. (2023) for more rigorous evaluation with better test coverage. As shown in Table 12, with HLP applied to FIM data only (i.e., ), the performance on L2R tasks sometimes shows a slight degradation. We hypothesize that applying HLP to only causes unbalanced training on and parts.

### v4

@@ Abstract
Yifeng Ding1††thanks: Work done at AWS AI Labs. Hantian Ding2 Shiqi Wang3∗ Qing Sun2 Varun Kumar2 Zijian Wang3∗ 1University of Illinois Urbana-Champaign 2AWS AI Labs 3Meta yifeng6@illinois.edu {dhantian,qinsun,kuvrun}@amazon.com {tcwangshiqi,zijianwang}@meta.com
@@ 1 Introduction
Fill-in-the-Middle (FIM), or infilling, has become essential for modern code development, where programmers frequently need to insert or modify code between existing sections rather than writing linearly from start to end (openaifim; incoder). While large language models have shown remarkable capabilities in code generation, the FIM task poses unique challenges that go beyond traditional left-to-right generation. A model must not only generate code that follows from the preceding context (), but also smoothly connect to the subsequent code () – a task that requires careful planning and foresight.
Current approaches to FIM typically reorder the input sequence and apply standard next-token prediction (NTP) training (starcodertwo; dscoder; dscodervtwo; qwentwofive). However, as illustrated in Figure 1, this methodology has a fundamental limitation: models struggle to maintain coherence over longer sequences and often fail to create smooth transitions to the right context (openaifim). While existing benchmarks attempt to address this through rule-based post-processing (e.g., truncating generated code based on line count (repoeval; repoformer) or program structure (crosscodeeval; safim)), such methods rely on dataset-specific assumptions that do not generalize to real-world scenarios where both left and right contexts can be arbitrary.
@@
| RepoEval (repoeval) | Truncate generation to the same number of lines as in ground truth. |
| CrossCodeLongEval (repoformer) | |
| CrossCodeEval (crosscodeeval) | Truncate generation at the first complete statement. |
| SAFIM (safim) | Stop when the target program structure in ground truth is generated. |
@@ 2 Post-processing for Fill-in-the-Middle
Most existing FIM works rely on post-processing to truncate code completions generated by LLMs for infilling tasks (safim; repoeval; crosscodeeval; repoformer). While such post-processing can enhance the FIM performance, we argue that they fundamentally depend on dataset-specific assumptions that make them impractical for real-world scenarios (§2.1). Through evaluation, we show that the performance of FIM existing code models drops significantly without post-processing, suggesting that post-processing conceals the fundamental struggles of models with code filling (§2.2). Furthermore, we show that this limitation stems from models’ inability to plan coherent completions given a fixed – a challenge that post-processing alone cannot address.
@@ 3 Horizon-Length Prediction
Next-token prediction (NTP) training is conducted on the document , which aims to minimize the following cross-entropy loss (where refers to the LLMs being trained):
@@ 4 Experiments
| + HLP | 49.4 | 54.5 | 61.8 | 65.8 | 57.9∗ |
Training. We conduct continual pre-training on a set of code LLMs of different model families and sizes to validate the effectiveness of HLP. Specifically, DeepSeek-Coder-Base 1.3B/6.7B (dscoder) and StarCoder2 3B/7B (starcodertwo) are involved in our experiments. We use AdamW (adamw) as the optimizer with and . We use a cosine learning rate scheduler with a peak learning rate equal to that at the pre-training end. All models are trained for 200K steps with a global batch size of 512.
Dataset. We use a subset of the-stack-v2 (starcodertwo) including Python, Java, C++, and C#. Following existing works (dscoder; starcodertwo), FIM rate is set to . We employ Best-fit Packing (ding2024fewertruncationsimprovelanguage) to group multiple files into each training sequence while masking out cross-file attention. The -- split is applied to each file independently rather than the whole training sequence.
@@ 4.1 Syntax-Aware Multilingual Code FIM
We use SAFIM (safim), a syntax-aware and multilingual code Fill-in-the-Middle benchmark, to evaluate the effectiveness of HLP. SAFIM focuses on syntax-aware completions of program structures, covering algorithmic block (i.e., Algo and Algo), control-flow expression (i.e., Control), and API function call (i.e., API). It consists of 17,720 examples from four different programming languages, including Python, Java, C++, and C#. SAFIM employs execution-based evaluation and reports pass@1 as the evaluation metric. As shown in Table 3, compared with NTP only, adding HLP achieves up to 5% improvements on average across all the studied code LLMs. Importantly, the improvement is consistent across languages and program structures, demonstrating HLP’s ability to enhance FIM capabilities regardless of language or completion context.
@@ 4.2 Repository-Level Cross-File Code FIM
In addition to single-file FIM evaluation with SAFIM, we also evaluate the effectiveness of HLP on repository-level code Fill-in-the-Middle in cross-file scenarios via CrossCodeEval (crosscodeeval), CrossCodeLongEval (repoformer), and RepoEval (repoeval). CrossCodeEval (Python) and CrossCodeLongEval are two repository-level cross-file benchmarks that leverage more than 1500 raw Python repositories to construct 12,665 examples across line, chunk, and function completion tasks, which are used for a more rigorous evaluation. RepoEval is another repository-level cross-file code completion benchmark consisting of 3,655 line, API, and function completion tasks created from 32 Python repositories. We follow existing work (repoformer) to evaluate the model’s FIM performance on these benchmarks and use Exact Match (EM) and Edit Similarity (ES) as our evaluation metrics. As shown in Table 4, adding HLP provides consistent improvements for all models across different benchmarks and completion tasks. Specifically, HLP achieves up to 24% improvements on EM and 9% improvements on ES relatively, showing its significant effectiveness.
@@ 4.3 Code Repair via Fill-in-the-Middle
To assess HLP’s impact on practical applications beyond code completion, we evaluate the performance of code repair using Defects4J (defects). Defects4J consists of open-source bugs found across 15 Java repositories. Following existing works (fitrepair; chatrepair), we collect 313 single-hunk bugs from Defects4J that can be fixed by replacing or adding a continuous code hunk. Specifically, for each bug, models are prompted to generate the correct code hunk (i.e., patch) given the left and right contexts of the buggy code hunk, with correctness verified by project test suites. As shown in the “Code Repair” section of Table 5, adding HLP during training results in relatively up to 18% more bugs fixed by the model222Note that Defects4J is a small dataset with only 313 examples and models can only solve 30-60 out of those, which makes it hard to obtain statistically significant differences with greedy decoding., suggesting that enhanced planning capabilities translate directly to better bug-fixing performance.
@@ 4.4 Code Reasoning via Fill-in-the-Middle
We further examine whether HLP’s planning benefits extend to code reasoning tasks. We use CRUXEval (cruxeval) which comprises 800 Python functions paired with two distinct tasks: CRUXEval-I, where LLMs need to predict the input from the known output, and CRUXEval-O, where LLMs are required to predict the output based on the given input.
@@ 5.1 NTP Alone Cannot Yield Horizon Awareness
To quantify this, we design a probing task that attempts to predict the remaining completion length from models’ hidden states. We extract hidden states from 20K code snippets (approximately 7.8M tokens) using models trained with and without HLP and split the data to ensure no sequence-level overlap between train and test. Using these hidden states as input and normalized remaining token counts as targets, we fit linear regression models while keeping the underlying transformer parameters frozen.
@@ 6 Related Work
Fill-in-the-Middle for Code Language Models Large language models trained on massive source code data have demonstrated great potential in various applications for software development. While early models such as Codex (humaneval) and CodeGen (codegen) only support Left-to-Right (L2R) generation, Fill-in-the-Middle (or infilling) has attracted increased attention because right context naturally carries an indispensable part of information for completing code in the middle (incoder; openaifim). Subsequently, FIM training has become a common practice widely adopted by most code LLMs, such as StarCoder (starcoder; starcodertwo), DeepSeek-Coder (dscoder; dscodervtwo), and Code Llama (codellama).
@@
An alternative approach is to train two language models in different directions, with one generating from left to right and the other from right to left, and have the two generations meet in the middle (meet-in-the-middle). Nevertheless, the L2R model does not have access to the right context, and vice versa, which impedes holistic planning that takes into account the context from both sides.
Planning and Lookahead in Language Generation Standard decoder-only models are trained with next-token prediction and used to sequentially predict one token at a time, conditioned only on past tokens, in an auto-regressive manner. One drawback of this paradigm is that models are not aware of future tokens during decoding. The token that maximizes the conditional probability at current step may lead to suboptimal continuation, and consequently the model can fail to compose a fluent and sensible generation that meets human requirements. Various decoding techniques have been proposed to address the problem through tree search with lookahead heuristics, particularly for constrained generation problems (NeuroLogic; DeAL). While these methods are training-free, they inevitably incur additional cost of inference complexity.
Apart from those, multitoken proposed to predict multiple tokens from a single hidden state during both training and inference, which was shown to achieve stronger performance on coding tasks with no computation overhead. While multi-token prediction enhances models’ planning capability within the tokens predicted together (), we argue that with a small , the limited horizon is usually insufficient for planning in the case of infilling as the connection from to only happens towards the end of the generation. In contrast, HLP adopts a global and arbitrary long horizon over all future tokens by counting the remaining generation budget, which more effectively helps models to close the generation fluently with early planning444Please refer to Appendix §A.3 for more details..
@@ A.3 Comparing HLP with Multi-token Prediction
| + HLP | 41.3 | 46.1 | 53.4 | 59.0 | 50.0 |
| + multi-token prediction | 38.1 | 41.3 | 51.7 | 55.8 | 46.7 |
As discussed in §6, we argue that multi-token prediction (multitoken) is insufficient for planning in FIM, because multi-token prediction only enhances models’ planning capability for a short and limited horizon, which does not suites FIM well as the connection from to happens over a long horizon. Instead, HLP focuses on long-horizon planning and is more effective for FIM. We conduct an experiment following the same settings in §4 to compare HLP with multi-token prediction on DeepSeek-Coder-Base 1.3B by predicting the next 4 tokens (multitoken). We report their performance on SAFIM. As shown in Table 10, while adding HLP to NTP largely improves the model’s performance on SAFIM, multi-token prediction fails to do so. These results provide empirical evidence that long-horizon planning capabilities brought by HLP is essential for advancing FIM performance.
@@ A.4 HLP Improves Self-Infilling Performance
Recently, some strategies have been proposed to conduct L2R code generation based on FIM capabilities and self-infilling (zheng2024selfinfillingcodegeneration) is one of the representative strategies of this kind. Given an initial input prompt , self-infilling performs L2R code generation by using FIM decoding to (1) generate based on and then (2) generate based on and previously generated . Furthermore, self-infilling proposes a looping mechanism to improve the generated code iteratively, where it first uses L2R decoding to generate a new ′ based on and previously generated and then uses FIM decoding to generate a new ′ based on and this newly generated ′. This looping procedure can be continued for multiple rounds to obtain greater improvements.
@@
| + HLP | 30.5 | 31.7 | 31.7 | 32.3 | 32.9 |
Since self-infilling relies on FIM capabilities for L2R code generation, it is interesting to study whether HLP can also enhance its performance. To this end, we evaluate the self-infilling performance of DeepSeek-Coder-Base 1.3B trained with and without HLP on HumanEval using greedy decoding, following the setting of the original paper (zheng2024selfinfillingcodegeneration). We report models’ performance on HumanEval with ranging from 0 to 4, where denotes the number of times the decoding process goes through the loop and represents that the looping mechanism is not activated. As shown in Table 11, the model trained with HLP consistently outperforms the model trained without HLP across different number of self-improving steps. Furthermore, with more self-improving steps, while the performance of the model trained without HLP gets stuck, the performance of the model trained with HLP continues to showcase steady improvements.
@@ A.5 Effect of HLP on Left-to-Right Performance
While HLP have significantly improved the FIM performance of LLMs, we also study its impact on the L2R code completion. To this end, we evaluate L2R performance on HumanEval (humaneval) and MBPP (mbpp) with DeepSeek-Coder-Base 1.3B. We further employ HumanEval+ and MBPP+ from EvalPlus evalplus for more rigorous evaluation with better test coverage. As shown in Table 12, with HLP applied to FIM data only (i.e., ), the performance on L2R tasks sometimes shows a slight degradation. We hypothesize that applying HLP to only causes unbalanced training on and parts.